OpenStudy (anonymous):

281 = 39 - u -8 = y +7 47 - y = 193 v - 5 = -4 please help me!!!

OpenStudy (anonymous):

281 = 39 - u 281 - 39 = -u 242 = -u u = -242

OpenStudy (anonymous):

the others are similar

OpenStudy (anonymous):

um for the 47 - y = 193 would that be -240??

OpenStudy (anonymous):

240 is too big 47 - - 240 is larger than 193

OpenStudy (anonymous):

sorry i mis typed -146?

OpenStudy (anonymous):

that seems about right

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

47+146 (double negative makes positive) = 193 so that checks out indeed

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):